OpenStudy (anonymous):

10. w – 2 = –3 (1 point) w = –3/2 w = –5 w = –1 w = 6 11. 2.8 = 2y (1 point)y = 1.4 y =1/2 y = 2.8 y = 0.6

OpenStudy (anonymous):

12. –9 =r/4 (1 point )r = 36 r = 13 r =9/4 r = –36

OpenStudy (mathstudent55):

For 10. Add 2 to both sides. For 11. Divide both sides by 2. For 12. Multiply both sides by 4.

OpenStudy (anonymous):

14. Which equation is an identity? (1 point)11 – (2v + 3) = –2v – 8 5w + 8 – w = 6w – 2(w – 4) 7m – 2 = 8m + 4 – m 8y + 9 = 8y – 3